Grade 12Physical Chemistry

calculate the mole fraction of benzene in a solution containing 40% by mass of toulene.?

Shriya Mehrotra
11 Years agoGrade 12

4 Answers

Sunil Kumar FP
11 Years ago
let the mass of the total mixture be 100gm
mass of the benzene=40gm
mass of the toluene =60gm
molecular weight of benzene and toluene =78 and 92
mole fraction of benzene =mole of benzene/total mole
=40/78/(40/78+60/92)
=.44
Shriya Mehrotra
11 Years ago
sr but he ans is 0.638
Sunil Kumar FP
11 Years ago
Yes you are correct.I made the caculation using the total mass of the mixture is 100gm
According to the question
if the mass of touene is 100 gm
mass of benzene=40 gm
moles of toluene=92/100 =100/92 =1.08
moe of benzene=40/78 =.51
moe fraction of benzene=.51/(1.08+.51)
=.319
please check the answer and let me know.I think the answer is wrong
britto george
11 Years ago
this is a simple question and you can get good things from solving this so just try